Show that for two scalar fields f and g:
∇.[∇f ×( f ∇g)]= 0
1
Expert's answer
2019-03-19T11:20:30-0400

This is most easily proved by using the Kronecker totally anti-symmetric symbol "\epsilon_{ijk}" such that "\epsilon_{123}=0". Given two vectors "a" and "b" with the Cartesian components, respectively, "a_i" and "b_i", the Cartesian components of their vector product are given by

"\left[ a \times b \right]_i = \sum_{jk} \epsilon_{ijk} a_j b_k \, ,"

and their scalar product is "a \cdot b = \sum_i a_i b_i". Applying these formulas, and using the simple relation "f \nabla f = \frac12 \nabla f^2", we have

"\nabla \cdot \left[ \nabla f \times \left( f \nabla g \right) \right] = \nabla \cdot \left[ f \nabla f \times \nabla g \right] = \frac12 \nabla \cdot \left[ \nabla f^2 \times \nabla g \right] =""\frac12 \sum_i \partial_i \left[ \nabla f^2 \times \nabla g \right]_i = \frac12 \sum_{ijk} \partial_i \left( \epsilon_{ijk} \partial_j f^2 \partial_k g \right) =""\frac12 \sum_{ijk} \epsilon_{ijk} \left( \partial_i \partial_j f^2 \partial_k g + \partial_j f^2 \partial_i \partial_k g \right) = 0 \, ."


The last equality is valid because "\partial_i \partial_j f^2" is symmetric with respect to {ij}, and "\partial_i \partial_k g" is symmetric with respect to {ik}, whereas "\epsilon_{ijk}" is totally anti-symmetric.